OpenStudy (anonymous):

Find all degree solutions: 2sin^2 3Ө + 3sin3Ө+1=0

OpenStudy (turingtest):

let\[u=\sin\theta\]then try to factor it

OpenStudy (anonymous):

\[2\sin^2 3Ө + 3\sin3Ө+1=0\] let \[\sin3 \theta = x\]

OpenStudy (turingtest):

^^^ sorry, that's what I meant ;)

OpenStudy (anonymous):

It would give 2x^2 + 3x +1=0

OpenStudy (anonymous):

so let x=sin3theta right ?

OpenStudy (anonymous):

so after that what do i do ?

OpenStudy (anonymous):

theta=pi/2,7pi/6,11pi/6,..... and5pi/12,13pi/12,....

OpenStudy (anonymous):

please explain it in degrees thanks :)

OpenStudy (anonymous):

2x^2 + 2x +x +1 2x(x+1) +1(x+1) (2x+1)(x+1) = 0 x=1/2 or x= -1

OpenStudy (anonymous):

Then sin 3theta = x sin3 theta= 0.5 theta = sin3^-1 times 0.5 which equals 9.55366

OpenStudy (anonymous):

i just explaineed

OpenStudy (anonymous):

now for all the answers, you first need to subtract from 180 degrees that answer, and add to 180 degrees that answer, and subtract from 360 degrees that answer.

OpenStudy (anonymous):

wait what? all of that for multiple answers ? how do i know which one to add 360 and which not?

OpenStudy (anonymous):

can't i use quadratic formula to get x ?

OpenStudy (anonymous):

yes, but you have one answer, and then you need to subtract it from different degrees to get all the answers.

OpenStudy (anonymous):

is 30 and 90 are some of the answers ?

OpenStudy (anonymous):

90 and 75 are 1 of the anwer

OpenStudy (anonymous):

I'm not sure If we have multiples, such as cos(2x), then we also multiply up the possible number of solutions. In "general", as a rule of thumb, expect the following, within any 360 degree range: sin(x) --> 2 solutions sin(2x) --> 4 solutions sin(3x) --> 6 solutions sin(4x) --> 8 solutions sin(5x) --> 10 solutions etc http://www.algebra.com/algebra/homework/Trigonometry-basics/solving-trigonometric-equations.lesson

OpenStudy (anonymous):

so in this case it could have up to 6 solutions ?

OpenStudy (anonymous):

Yes. Just check more sites to see how to solve it for sin3x

OpenStudy (anonymous):

so did you found out all the answers ?

OpenStudy (anonymous):

if i found 90 is one of the answer do i take 180 and subtract it or something?

OpenStudy (anonymous):

No, I forgot it wasn't sin2x until I went to this site... so I forgot the range you need to subtract and add to.. I'm sure sahii knows.

OpenStudy (anonymous):

@sahil51993 come here :D

OpenStudy (anonymous):

i used quadratic and i got x=-.5 and x=-1 =.=

OpenStudy (anonymous):

we got sin3x =-1 or -1/2 now we know that sinx is -1 for 3pi/2 and so on +2pi and sinx is -1/2 when angle is 5pi/4 and so on hence we can equate these to 3x and find the requied value of x in terms of pi and later put pi=180 degrees to calculate the answer

OpenStudy (anonymous):

@sahil51993 so we put -1=3x and -1/2=3x just solve for x ?

OpenStudy (anonymous):

no we put sin 3x=-1 and sin 3x=-1/2

OpenStudy (anonymous):

@sahil51993 after that can't i just use Sin inverse to get the answers?

OpenStudy (anonymous):

sin^-1(-1/3) and sin^-1(-1/6) ?

OpenStudy (anonymous):

@sahil51993 i got -19.47 and -9.5

OpenStudy (anonymous):

yes u can but do take all the possibilities

OpenStudy (anonymous):

@sahil51993 so do you take sin^-1(1/3) and got 19.47 and take that 19.47+180/3 to get answer or what? please help me pleaseee

OpenStudy (anonymous):

sorry for keep bothering you, but i really need help >.<

OpenStudy (anonymous):

nopes 1/3 sin^-1 (-1) gives 1 value as 270/3 as sin^-1(-1) is 270 and so on 1/3 sin^-1 (-1/2) gives 1 value as 225/3 as sin^-1(-1/2) is 225 and so on

OpenStudy (anonymous):

oh you gotta put 1/3 in front ok

OpenStudy (anonymous):

wait 1/3sin^1(-1) does not give you 270/3 wth

OpenStudy (anonymous):

i got -30 from that

OpenStudy (anonymous):

it does sinx=-1 for 270,630,990,.......

OpenStudy (anonymous):

no....my calc, oh wait you don't put it in calc ?

OpenStudy (anonymous):

no i can calculate it mentally

OpenStudy (anonymous):

i tried it in the calc but it does not give me that

OpenStudy (anonymous):

I got 1 thing wrong sin ^-1(-1/2)=210,570,930,.....

OpenStudy (anonymous):

that give me -30 =.=

OpenStudy (anonymous):

i dont understand what im doing wrong :|

OpenStudy (anonymous):

u can say that because sinx is an odd function so sin-30=-sin30 so yes but this seems incorrect in practical life where u cannot represent answer in negative vale actually ur question is not proper it should as us value of theta in a particular range......get it....and i have my exam day after tomorrow so please make it quickly

OpenStudy (anonymous):

oh sorry, ok just one more thing, please explain how you got from 270 to 630 to 990 what did you do ?

OpenStudy (anonymous):

sinx=-1 for 270 and again when it completes a peroid and comes back to -1 value is increased by 360 and so on

OpenStudy (anonymous):

ok i got it thanks!!